OpenStudy (anonymous):

Select all of the following that are ordered pairs of the given function. f(x) = 3 - 2x (-2, -1) (-1, 5) (0, 3) (1, 0) (2, -1)

OpenStudy (anonymous):

The points are in the format : \((x, f(x))\), so first value is \(x\) and other value is \(f(x)\)..

OpenStudy (anonymous):

One by one, put \(x\) and \(f(x)\)value from the options in the expression given to you, and check whether you get the equation satisfied or not..

OpenStudy (anonymous):

Like I try your first option: There \(x = -2\) and \(f(x) = -1\), put these values in the expression given: \[-1 = 3- 2(-2) \\ -1 = 3 + 4 \\ -1 \ne 7\]

OpenStudy (anonymous):

so three and five are correct?

OpenStudy (anonymous):

So, option 1 is not satisfying the equation..

OpenStudy (anonymous):

Yes, Excellent...

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

\(\dagger\)